A chemist has 40 ounces of a solution of iodine and alcohol, which is 15% iodine. How much pure iodine must be added to make a solution that is 20% iodine?
Answer by mananth(16949)

--------percent ---------------- quantity
Iodine 15 ---------------- 40 oz
Iodine 100 ---------------- x oz
Mixture 20 ---------------- 40 + x oz

15*40+100x=20(40+x)
600+100x=800+20x
100x-20x=800-600
80x=200
/ 80
x = 2.5
2.5 oz of Iodine